OpenStudy (anonymous):

How to convert 1-1/(jw+1)^2 to a real value and an angle?

OpenStudy (anonymous):

j being complex

OpenStudy (anonymous):

w=omega

OpenStudy (ash2326):

\[1-\frac{1}{(1+jw)^2}\] Let's simplify this first \[\frac{(1+jw)^2-1}{(1+jw)^2}\] We will get \[\frac{-w^2-2jw}{1-w^2+2jw}\] DO you get this part ?

OpenStudy (anonymous):

I see how you did it but do we need it in that form?

OpenStudy (ash2326):

yes, it's easy that way

OpenStudy (ash2326):

Let's find the real value first ( or absolute value)

OpenStudy (anonymous):

Okay I'm just having trouble like figuring how you find it from the form j

OpenStudy (ash2326):

\[\frac{|-w^2-2jw|}{|1-w^2+2jw|} \] This is the real value\[\frac{\sqrt{(-w^2)^2+(-2w)^2}}{\sqrt{(1-w^2)^2+(2w)^2}}\]

OpenStudy (ash2326):

a+jb Real value \[\sqrt{a^2+b^2}\]

OpenStudy (anonymous):

oh you do it for both top and bottom

OpenStudy (anonymous):

oh I see so you try to get it in the form a+jb

OpenStudy (ash2326):

yes :)

OpenStudy (ash2326):

Yes, we can do it both for numerator and denominator

OpenStudy (anonymous):

oh sweet okay how do you get the angle? is there some diagram we can see

OpenStudy (anonymous):

like which is x and which is y

OpenStudy (ash2326):

Yes, we can find the angle separately

OpenStudy (ash2326):

\[\frac{-w^2-2jw}{1-w^2+2jw}\] Angle get subtracted from numerator and denominator \[{\tan^{-1} (\frac{-2w}{-w^2})}-{\tan^{-1}(\frac{2w}{1-w^2})}\] Do you get this part ?

OpenStudy (anonymous):

are you just dividing b/a?

OpenStudy (ash2326):

Yes, but one more \[a+ib=> \tan^{-1} \frac b a\] only if a>0 and b>0 and a<0, b<0 if a>0,b<0 or a<0 b>0 =>\(180-\tan^{-1}{\frac ba }\)

OpenStudy (ash2326):

There is a catch here, if w^2 <1 then it'll be \[{\tan^{-1} (\frac{2w}{w^2})}-{\tan^{-1}(\frac{2w}{1-w^2})}\] if w^2>1 then \[{\tan^{-1} (\frac{2w}{w^2})}-(180-{\tan^{-1}(\frac{2w}{w^2-1}))}\]

OpenStudy (anonymous):

Oh whoa serious.. I am not given any info on w

OpenStudy (anonymous):

but it's alright I'm going to assume it's greater one question though

OpenStudy (anonymous):

For the real part though I am looking at my text and when trying your method for the real part it doesn't seem to work? They got 4.47 but if you do it your way is it not 4?

OpenStudy (ash2326):

cool :)

OpenStudy (ash2326):

it's easy \[\frac{10\times \sqrt{2^2+4^2}}{\sqrt{6^2+8^2}}=10\times \sqrt {20}/10=\sqrt {20}\] Did you also do the same thing?

OpenStudy (anonymous):

oh you can't multiply the 10 into the brackets? to get 20 + j40?

OpenStudy (ash2326):

You can But it'll come out eventually so I kept it outside

OpenStudy (anonymous):

oh but if you do it with 20+j40 can't you just assume a=20 and b =40?

OpenStudy (anonymous):

ah sorry for all the questions

OpenStudy (ash2326):

you can, it's right No worries. I'm here to help you questions are good :)

OpenStudy (anonymous):

lol okay but if you do then you get sqrt of 1600 which is 40 / 10?

OpenStudy (anonymous):

which gives you 4

OpenStudy (ash2326):

\[20^2+40^2=400+1600=2000\]

OpenStudy (anonymous):

Lol omg I'm so dumb so sorry for the dumb questions

OpenStudy (anonymous):

yeah it works out haha thanks man

OpenStudy (ash2326):

No problem :)

OpenStudy (anonymous):

yeah seriously thanks for your help really got it now!!

OpenStudy (ash2326):

Glad to help you :)

OpenStudy (anonymous):

have a good night!

OpenStudy (ash2326):

Night